Michael has exactly 100 coins,all nickels and dimes. She counted her money and it came to $8.10. How many of each coin does michael have?
Answer by MRperkins(300)

n=nickels
d=dimes
n+d=100 <====if you add the number of nickels (n) and dimes (d) you get 100
.05n+.10d=8.10 <=====a nickel is worth .05 and a dime is worth .10.
n=100-d <=====You can either use substitution or elimination. Substitution is the most basic method.
.05(100-d)+.10d=8.10
5-.05d+.10d=8.10
5+.05d=8.10
-.05d=-3.1
d=62
n=100-62 or n=38
therefore, there are 62 dimes and 38 nickels.
.
Do you want your math questions answered quickly and completely? Do you want to ask follow up questions like, “How do you do that?” or “Why did you do that?”? Email your questions to justin.sheppard.tech@hotmail.com and I will answer your questions within 12 hours of submission (usually within 2-3 hours of submission). I charge $2 per question and I will provide a .pdf file of the problem worked out. I will also offer to discuss the problem with you over the phone, through skype, or via an online whiteboard. Payments to be made securely through PayPal AFTER you are satisfied with the answer and services that I provide. This eliminates the possibility of being scammed (worst case scenario is that I provide you with your answer and you don’t pay me).
.
Respectfully,
Justin
justin.sheppard.tech@hotmail.com